Find the number of sides of the regular polygon whose interior angles each measure 156 degrees
Found 2 solutions by Alan3354, oberobic:
Answer by Alan3354(31538)

Find the number of sides of the regular polygon whose interior angles each measure 156 degrees
------
Ext angle = 180 - Int angle = 24
n = 360/24 = 15 sides

Answer by oberobic(2304)

Total interior angles = (n-2)*180
.
So, the average interior angle = ( (n-2)*180 ) / n
.
((n-2)*180)/n = 156
.
(n-2)*180 = 156n
180n -360 = 156n
24n = 360
n = 15
.
The polygon has 15 sides.